Machines A and B make components, which are then placed on a conveyor belt. Of those made by machine A, 93% are acceptable. Of those made by machine B, 95% are acceptable. Machine A makes 60% of the components and machine B makes the rest. Find the probability that a component selected at random from the conveyor belt is a) made by machine A b) made by machine A and acceptable c) made by machine B and acceptable d) made by machine ad uaeceptable e acceptable f) made by machine A given it is acceptable
Answer #1
A B total
accceptab;e 0.558 0.38 0.938
not acceptable 0.042 0.02 0.062
total 0.6 0.4 1

a) P(machine A) = 0.60  
  
b) P(machine A and acceptable) = 0.60*0.93 = 0.558  
  
c) P(machine B and acceptable) = 0.60*0.95 = 0.38  
  
d) P(machine and unaccepetable) = 0.60*(1-0.95) = 0.02  
  
e) P(acceptable) = 0.558+0.38 = 0.938  
  
f) P(machine A | acceptable) = P(machine A and acceptable)/P(accpetable) = 0.558/0.938= 0.595